Don randomly draws two cards from a standard deck of 52 cards. He does not replace the first card. What is the probability that both cards are aces?
Answer by jim_thompson5910(35256)

P(1st card is ace) = 4/52 = 1/13

P(2nd card is ace) = 3/51

-------------------------------------------------------

P(both cards are aces) = P(1st card is ace)*P(2nd card is ace)

P(both cards are aces) = (1/13)*(3/51)

P(both cards are aces) = 3/663

P(both cards are aces) = 1/221

------------------------------------------------------------------------------------------------------------------------

If you need more one-on-one help, email me at jim_thompson5910@hotmail.com. You can ask me a few more questions for free, but afterwards, I would charge you ($2 a problem to have steps shown or $1 a problem for answer only).

Alternatively, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Any amount is greatly appreciated as it helps me a lot. This donation is to support free tutoring. Thank you.

Jim
------------------------------------------------------------------------------------------------------------------------